# The Wavefront Set: Bounds for the Langlands parameter

### v1

@@ 1.1. Langlands parameters
We briefly recall the conjectural Langlands parametrization of the irreducible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representations, in the formulation of [Vo], generalizing the Kazhdan-Lusztig geometrical parametrization of representations with Iwahori-fixed vectors. A Langlands parameter is a group homomorphism
@@
Denote by $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})=\Pi(\varphi)$ the corresponding L-packet, whose members should be parametrized by the local systems $\mathcal{L}^{\vee}$ on ${{\mathcal{O}}}^{\vee}$. A desideratum is that $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$ consists of tempered representations if and only if ${{\mathcal{O}}}^{\vee}$ is dense in ${\mathfrak{h}}^{\vee}_{q}$.
@@ 1.2. Aubert-Zelevinsky involution
Let ${\mathsf{AZ}}$ denote the normalized Aubert-Zelevinsky involution [Au] on $\mathsf{Irr}{\mathbf{G}}^{\omega}({\mathsf{k}})$ for every form ${\mathbf{G}}^{\omega}({\mathsf{k}})$. If $\varphi_{\pi}$ is the Langlands parameter of $\pi$, then it is expected that $\varphi_{\pi}(W_{k})=\varphi_{{\mathsf{AZ}}(\pi)}(W_{k})$. Moreover, if $\pi=\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$, it is expected that the Langlands parameter of ${\mathsf{AZ}}(\pi)$ is related to that of $\pi$ via the Fourier-Deligne transform on the appropriate space of perverse sheaves of Langlands parameters. This is known to hold for all representations with Iwahori-fixed vectors by [EM], and more generally it is expected to hold for unipotent representations in the sense of Lusztig.
@@
Let $\operatorname{FT}$ denote the Fourier-Deligne transform from $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q}$ to $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q^{-1}}$. Let $\mathsf{IC}({{\mathcal{O}}},{{\mathcal{L}}})$ denote the simple perverse sheaf given by the intersection cohomology complex defined by the equivariant irreducible local system ${{\mathcal{L}}}$ supported on ${{\mathcal{O}}}$. Then $\operatorname{FT}(\mathsf{IC}({{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))=\mathsf{IC}({{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}})$ for an orbit ${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$ in $\mathfrak{g}^{\vee}_{q^{-1}}$ and an equivariant local system ${{\mathcal{L}}}^{\vee}_{\mathsf{FT}}$. If $\mathfrak{g}^{\vee}=\mathfrak{gl}(n)$, so that all local systems that appear are trivial, [EM, Proposition 7.2] shows that
@@
where *** in the left hand side denotes the contragredient. In our conjectures below, we will be interested in the nilpotent part of the Langlands parameter of ${\mathsf{AZ}}(\pi)$, that is in $\mathbb{O}^{\vee}_{\operatorname{FT}}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$
@@ 1.4. Unramified wavefront set
If $\pi$ is an admissible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation, let ${}^{K}\!\mathsf{WF}(\pi)$ denote the unramified wavefront set of $\pi$ as in [Ok]. This is a (finite) collection of nilpotent ${\mathbf{G}}^{\omega}(K)$-orbits in $\mathbf{\mathfrak{g}}^{\omega}(K)$. Using the parametrization in loc. cit., there is a one-to-one correspondence between the set of nilpotent ${\mathbf{G}}^{\omega}(K)$-orbits and pairs [^1] Here we need split over unramified extension.
$({{\mathcal{O}}},C):\ {{\mathcal{O}}}\text{ nilpotent $G$-orbit in }\mathfrak{g},\ C\text{ conjugacy class in }A({{\mathcal{O}}}),$
where $A({{\mathcal{O}}})$ is a group of components of the centralizer $C_{G}(u)$ in $G$ of a representative $u$ of ${{\mathcal{O}}}$. By abuse of notation, we think of ${}^{K}\!\mathsf{WF}(\pi)$ as a collection of such pairs $({{\mathcal{O}}},C)$. For each ${{\mathcal{O}}}$ choose a representative $u\in{{\mathcal{O}}}$. For each $C$, we may choose a semisimple representative $\tau$ in $C_{G}(u)$. Write
@@
this is independent of the choice of $u$ and $\tau$. Also write
@@ 1.5. Conjectures
${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=d^{\vee}(\mathbb{O}^{\vee}),$ (1.4)
where ${\mathbb{O}}^{\vee}$ is the $G^{\vee}$-saturation of the nilpotent Langlands parameter of $\pi$.
To exemplify Conjecture 1.1, we will verify it for the depth-zero supercuspidal representations of classical groups, using the Langlands parametrization of Lust-Stevens [LS], and of $G_{2}$, using Aubert-Xu [AX].
@@
If the representation $\pi(s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ is generic, then its geometric wavefront set is the principal orbit. For the conjecture to hold, we would need $\mathbb{O}^{\vee}_{\mathsf{FT}}=0$, so ${{\mathcal{O}}}^{\vee}$ would need to be the open dense orbit in ${\mathfrak{h}}^{\vee}_{q}$ and, at least when $H^{\vee}(s)$ is connected, ${{\mathcal{L}}}^{\vee}$ would have to be the trivial local system.
@@
Notice that Conjecture 1.6 is vacuously true for the nonsingular supercuspidal representations, in the sense of Kaletha [Kal], because ${\mathbb{O}}^{\vee}_{s}=0$ so $d^{\vee}({\mathbb{O}}^{\vee}_{s})$ is the principal geometric orbit. In this case, the upper bound is attained only for the generic constituents of the L-packets. In section 3.1, we will compute the wavefront sets of the regular depth-zero supercuspidals parameterized by DeBacker-Reeder [DBR].
@@
Suppose ${\mathbf{G}}({\mathsf{k}})$ is a symplectic, special orthogonal, or unitary group. Conjecture 1.6 holds for all depth-zero supercuspidal representations, using the Langlands parametrization of Lust-Stevens [LS].
We mention that Conjectures 1.1 and 1.6 are compatible with a conjecture about Arthur packets formulated by D. Jiang. Suppose $\varphi$ is a tempered L-parameter. Then ${\mathsf{AZ}}(\Pi(\varphi))$ is expected to be an Arthur packet. Let $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$ be one of these “anti-tempered” representations. Then $\pi^{\prime}={\mathsf{AZ}}(\lambda,s,{{\mathcal{O}}}^{\vee}_{s},{{\mathcal{L}}}^{\vee})$ for some local system ${{\mathcal{L}}}^{\vee}$. Conjecture 1.1 then says that we should have
@@
Acknowledgements. This paper builds on the previous work of the first-named author with Lucas Mason-Brown and Emile Okada. It is a pleasure to thank them for many useful conversations about the wavefront set. We also thank Baiying Liu and Freydoon Shahidi for sharing their work with us and for an interesting correspondence. D.C. was partially supported by the EPSRC grant “New Horizons” EP/V046713/1, 1/2021-12/2023.
@@ 2. Parabolic induction
In this section, we discuss the compatibility of our conjectures with parabolic induction. Let $\mathbf{P}=\mathbf{M}\mathbf{N}$ be a $\mathsf{k}$-rational parabolic subgroup of ${\mathbf{G}}$ with a $\mathsf{k}$-rational Levi subgroup $M$.
If $\pi=\operatorname{Ind}_{\mathbf{P}({\mathsf{k}})}^{{\mathbf{G}}({\mathsf{k}})}(\pi_{M})$ is a parabolically induced representation, where $\pi_{M}$ is an admissible $\mathbf{M}(\mathsf{k})$-representation, then the relation between the wavefront sets is well known, see Waldspurger [Wa]:
@@
Since the involution ${\mathsf{AZ}}$ commutes with parabolic induction [Au], i.e., in the Grothendieck group
@@ 3. Depth-zero supercuspidal representations
Lusztig gave an explicit description of $\mathsf{WF}(\mu)$, $\mu$ an irreducible representation of ${\mathbf{G}}^{\omega}_{x}({\mathbb{F}}_{q})$ in terms of his parametrization . Let $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$ be the dual finite reductive group. Then $\mu$ is parametrized by a triple $(\tau,u_{0}^{\vee},y)$, where $\tau$ is a semisimple element in $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$, $u_{0}^{\vee}$ is a special unipotent element in the centralizer $C_{({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})}(\tau)$ and $y$ is an element of the unipotent family $C_{u_{0}^{\vee}}$ parametrized by $u_{0}^{\vee}$. The element $y$ will not be important for us. Let $\widetilde{C}_{u_{0}^{\vee}}=C_{\widetilde{u}_{0}^{\vee}}$ be the family obtained by applying Alvis-Curtis duality to $C_{u_{0}^{\vee}}$. Let $d_{S}^{\vee}$ be the duality map in this finite field setting, this is defined exactly as above, see [Lu1, §13.3].
Theorem 3.2 ([Lu2]).
@@
${}^{K}\!\underline{\mathsf{WF}}(\pi(x,\sigma))=(d_{S}^{\vee}(\tau u_{0}^{\vee}),x),$ (3.1)
see [CMBO2]. Since we are in a depth-zero case, the Langlands parameter $\varphi$ is trivial on the wild inertia group, hence the image $\varphi(I_{\mathsf{k}})$ is a cyclic group generated by an element $\lambda\in{}^{L}\!G$ of finite order prime to $q$. Hence ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\lambda}$ is a pseudo-Levi subalgebra of $\mathfrak{g}^{\vee}$. Conjecture 1.6 is equivalent to
@@ 3.1. Regular supercuspidal representations
For all regular depth-zero supercuspidal representations [DBR], the conjecture is trivially true since in that case ${\mathfrak{h}}^{\vee}$ is just a Cartan subalgebra of $\mathfrak{g}^{\vee}$, so the only possible Langlands nilpotent element is 00. Hence the conjecture just says that ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq{{\mathcal{O}}}_{p}$, where ${{\mathcal{O}}}_{p}$ is the principal ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit, which is true, but trivial. On the other hand, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={{\mathcal{O}}}_{p}$ if and only if $\pi$ is a generic representation, and this is not always the case. In general, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ can be equal to the ${\mathbf{G}}(\bar{\mathsf{k}})$-saturation of the principal nilpotent orbit in a maximal pseudo-Levi, as we explain next.
@@
By abuse of notation, denote $x$ also the preimage in $X$ of an element in $X/(1-w\theta)X$. Let $\pi(\lambda,w,0,x)$ denote the supercuspidal representation attached to $(\lambda,w,x)$ in [DBR]. The 00 here stands for the zero nilpotent orbit, i.e., the image of the $SL(2)$ under the Langlands parameter must be trivial (since ${\mathfrak{g}^{\vee}}^{\lambda}=\mathfrak{t}^{\vee}$).
@@
The claim follows at once, after we recall the construction of $\pi(\lambda,w,0,x)$ from [DBR]. The element $x$ defines a cocycle $u_{x}\in H^{1}({\mathsf{k}},{\mathbf{G}})$ which gives the inner twist $\omega_{x}$ of ${\mathbf{G}}$ and $F_{x}$ a twisted Frobenius map and a maximal parahoric subgroup ${\mathbf{G}}_{x}$. Then ${\mathbf{G}}_{x}^{F_{x}}$ is a maximal parahoric subgroup of ${\mathbf{G}}^{\omega_{x}}({\mathsf{k}})$. The pair $(w,x)$ defines an $F_{x}$-minisotropic torus $T_{x}$ and $\lambda$ defines a depth-zero character $\chi_{x}$ of $T_{x}$. Then Deligne-Lusztig induction defines an irreducible cuspidal representation of the reductive quotient $\overline{{\mathbf{G}}_{x}^{F_{x}}}$
@@
4. Classical groups (after Lust-Stevens [LS])
Let ${\mathcal{G}}$ be a connected reductive group of classical type with connected center over ${\mathbb{F}}_{q}$, $q$ odd. Let $F$ be the Frobenius map. Let ${\mathcal{G}}^{*}$ be the dual group over ${\mathbb{F}}_{q}$ with dual Frobenius also denoted $F$. Let $\mathsf{Irr}{\mathcal{G}}^{F}$ be the set of equivalence classes of irreducible complex ${\mathcal{G}}^{F}$-representations. According to Lusztig’s classification, we can partition

### v2

@@ 1.1. Langlands parameters
We briefly recall the conjectural Langlands parametrization of the irreducible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representations, in the formulation of [Vo], generalizing the Kazhdan-Lusztig geometric parametrization of representations with Iwahori-fixed vectors. A Langlands parameter is a group homomorphism
@@
Denote by $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})=\Pi(\varphi)$ the corresponding L-packet, each of whose member $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ should be parametrized by the equivariant local systems $\mathcal{L}^{\vee}$ on ${{\mathcal{O}}}^{\vee}$. A desideratum is that $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$ consists of tempered representations if and only if ${{\mathcal{O}}}^{\vee}$ is dense in ${\mathfrak{h}}^{\vee}_{q}$.
@@ 1.2. Aubert-Zelevinsky involution
Let ${\mathsf{AZ}}$ denote the normalized Aubert-Zelevinsky involution [Au, Definition 1.5, Corollary 3.9] on $\mathsf{Irr}{\mathbf{G}}^{\omega}({\mathsf{k}})$ for every form ${\mathbf{G}}^{\omega}({\mathsf{k}})$. If $\varphi_{\pi}$ is the Langlands parameter of $\pi$, then it is expected that $\varphi_{\pi}(W_{\mathsf{k}})$ and $\varphi_{{\mathsf{AZ}}(\pi)}(W_{\mathsf{k}})$ are $G^{\vee}$ conjugate. Moreover, if $\pi=\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$, it is expected that the Langlands parameter of ${\mathsf{AZ}}(\pi)$ is related to that of $\pi$ via the Fourier transform [KS, §3.7] on the appropriate space of perverse sheaves of Langlands parameters, see also [EM, §3]. This is known to hold for all representations with Iwahori-fixed vectors by [EM, Theorem 0.1], and more generally it is expected to hold for unipotent representations in the sense of Lusztig.
@@
Let $\operatorname{FT}$ denote the Fourier transform from $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q}$ to $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q^{-1}}$, [EM, §3.2]. Let $\mathsf{IC}({{\mathcal{O}}},{{\mathcal{L}}})$ denote the simple perverse sheaf given by the intersection cohomology complex defined by the equivariant irreducible local system ${{\mathcal{L}}}$ supported on ${{\mathcal{O}}}$. Then $\operatorname{FT}(\mathsf{IC}({{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))=\mathsf{IC}({{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}})$ for an orbit ${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$ in $\mathfrak{g}^{\vee}_{q^{-1}}$ and an equivariant local system ${{\mathcal{L}}}^{\vee}_{\mathsf{FT}}$. If $\mathfrak{g}^{\vee}=\mathfrak{gl}(n)$, so that all local systems that appear are trivial, [EM, Proposition 7.2] shows that
@@
where * in the left hand side denotes the contragredient. In our conjectures below, we will be interested in the nilpotent part of the Langlands parameter of ${\mathsf{AZ}}(\pi)$, that is in $\mathbb{O}^{\vee}_{\operatorname{FT}}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$
@@ 1.4. Unramified wavefront set
If $\pi$ is an admissible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation, let ${}^{K}\!\mathsf{WF}(\pi)$ denote the unramified wavefront set of $\pi$ as in [Ok]. This is a (finite) collection of nilpotent ${\mathbf{G}}^{\omega}(K)$-orbits in $\mathbf{\mathfrak{g}}^{\omega}(K)$. Using the parametrization in loc. cit., there is a one-to-one correspondence between the set of nilpotent ${\mathbf{G}}^{\omega}(K)$-orbits and pairs
$({{\mathcal{O}}},C):\ {{\mathcal{O}}}\text{ nilpotent ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit in }\mathfrak{g}(\bar{\mathsf{k}}),\ C\text{ conjugacy class in }A({{\mathcal{O}}}),$
where $A({{\mathcal{O}}})$ is a group of components of the centralizer $C_{{\mathbf{G}}(\bar{\mathsf{k}}}(u)$ in ${\mathbf{G}}(\bar{\mathsf{k}})$ of a representative $u$ of ${{\mathcal{O}}}$. By abuse of notation, we think of ${}^{K}\!\mathsf{WF}(\pi)$ as a collection of such pairs $({{\mathcal{O}}},C)$. For each ${{\mathcal{O}}}$ choose a representative $u\in{{\mathcal{O}}}$. For each $C$, we may choose a semisimple representative $\tau$ in $C_{{\mathbf{G}}(\bar{\mathsf{k}})}(u)$. Choosing an isomorphism $\bar{k}\cong\mathbb{C}$, we may identify ${\mathbf{G}}(\bar{\mathsf{k}})$ with the complex group $G={\mathbf{G}}(\mathbb{C})$ with the same root datum, and think of the duality maps in section 1.3 as maps between ${\mathbf{G}}(\bar{\mathsf{k}})$ and $G^{\vee}$. Write
@@
which is independent of the choice of $u$ and $\tau$. Also write
@@ 1.5. Conjectures
${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=d^{\vee}(\mathbb{O}_{\pi}^{\vee}),$ (1.4)
where ${\mathbb{O}}_{\pi}^{\vee}$ is the $G^{\vee}$-saturation of the nilpotent Langlands parameter of $\pi$.
To exemplify Conjecture 1.1, we will verify it for the depth-zero supercuspidal representations of classical groups, using the Langlands parametrization of Lust-Stevens [LuSt], and of $G_{2}$, using Aubert-Xu [AX].
@@
If the representation $\pi(s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ is generic, then its geometric wavefront set is the principal orbit. For the conjecture to hold, we would need $\mathbb{O}^{\vee}_{\mathsf{FT}}=0$, so ${{\mathcal{O}}}^{\vee}$ would need to be the open dense orbit in ${\mathfrak{h}}^{\vee}_{q}$. Moreover, when $H^{\vee}(s)$ is connected, ${{\mathcal{L}}}^{\vee}$ would have to be the trivial local system.
@@
Notice that Conjecture 1.6 is vacuously true for the nonsingular supercuspidal representations, in the sense of Kaletha [Kal], because ${\mathbb{O}}^{\vee}_{s}=0$ so $d^{\vee}({\mathbb{O}}^{\vee}_{s})$ is the principal geometric orbit. In this case, the upper bound is attained only for the generic constituents of the L-packets. In section 3.1, we will compute the wavefront sets of the regular depth-zero supercuspidals parameterized by Kazhdan-Varshavsky [KV] and DeBacker-Reeder [DBR].
@@
Suppose ${\mathbf{G}}({\mathsf{k}})$ is a symplectic, special orthogonal, or unitary group. Conjecture 1.6 holds for all depth-zero supercuspidal representations, using the Langlands parametrization of Lust-Stevens [LuSt].
We mention that Conjectures 1.1 and 1.6 are compatible with a conjecture about Arthur packets formulated by D. Jiang. Suppose $\varphi$ is a tempered L-parameter. Then ${\mathsf{AZ}}(\Pi(\varphi))$ is expected to be an Arthur packet. Let $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$ be one of these “anti-tempered” representations. Then $\pi^{\prime}={\mathsf{AZ}}(\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{s},{{\mathcal{L}}}^{\vee}))$ for some local system ${{\mathcal{L}}}^{\vee}$. Conjecture 1.1 then says that we should have
@@
In section 5, we will verify Conjecture 1.9 for $G_{2}$, using the Langlands correspondences from [AX, GS1, GS2] and the wavefront set results of [JLS, LoSa].
@@
Remark 1.10.
We look at the two extreme cases of Conjectures 1.1 and 1.9.
  1. (a)
    Firstly, let us call an irreducible representation $\pi$ anti-generic if ${\mathsf{AZ}}(\pi)$ is generic. If $\pi$ is anti-generic, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$, the regular nilpotent orbit. The only way that the upper bound conjecture could hold then is if the Langlands nilpotent parameter of $\pi$ is 00. In other words, for every anti-generic representation the image of $\mathbb{C}\subset W_{\mathsf{k}}^{\prime}$ under the Langlands parameter is trivial.
    If $\pi$ is unipotent anti-generic, then it is known that it is spherical (with respect to the maximal special parahoric) and the desired condition on the Langlands parameter holds.
    If $\pi$ is supercuspidal, since ${\mathsf{AZ}}(\pi)=\pi$, the conjecture implies that every generic supercuspidal representation has trivial nilpotent Langlands parameter.
  2. (b)
    Secondly, suppose $\mathcal{O}^{\vee}=0$, so $d^{\vee}(G^{\vee}\cdot\mathcal{O}^{\vee})=\mathcal{O}_{\textup{reg}}.$ Then Conjecture 1.9(1) is trivially satisfied, but (2) says that there should exist a representation $\pi$ in the tempered L-packet such that ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$. In other words, every non-singular tempered L-packet should contain an anti-generic representation.
Acknowledgements. This paper builds on the previous work of the first-named author with Lucas Mason-Brown and Emile Okada. It is a pleasure to thank them for many useful conversations about the wavefront set. The second-named author would like to thank George Lusztig and Cheng-Chiang Tsai for insightful discussions. We also thank Baiying Liu and Freydoon Shahidi for sharing their work with us and for an interesting correspondence. D.C. was partially supported by the EPSRC grant “New Horizons” EP/V046713/1, 1/2021-12/2023.
@@ 2. Parabolic induction
In this section, we discuss the compatibility of our conjectures with parabolic induction. Let $\mathbf{P}=\mathbf{M}\mathbf{N}$ be a $\mathsf{k}$-rational parabolic subgroup of ${\mathbf{G}}$ with a $\mathsf{k}$-rational Levi subgroup $M=\mathbf{M}({\mathsf{k}})$.
If $\pi=\operatorname{Ind}_{\mathbf{P}({\mathsf{k}})}^{{\mathbf{G}}({\mathsf{k}})}(\pi_{M})$ is a parabolically induced representation, where $\pi_{M}$ is an admissible representation of $M$, then the relation between the wavefront sets is well known, see Waldspurger [Wa]:
@@
Since the involution ${\mathsf{AZ}}$ commutes with parabolic induction [Au, Theorem 1.7], i.e., in the Grothendieck group
@@ 3. Depth-zero supercuspidal representations
Lusztig gave an explicit description of $\mathsf{WF}(\mu)$, $\mu$ an irreducible representation of ${\mathbf{G}}^{\omega}_{x}({\mathbb{F}}_{q})$ in terms of his parametrization [Lu1, Theorem 4.23]. Let $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$ be the dual finite reductive group. Then $\mu$ is parametrized by a triple $(\tau,u_{0}^{\vee},y)$, where $\tau$ is a semisimple element in $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$, $u_{0}^{\vee}$ is a special unipotent element in the centralizer $C_{({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})}(\tau)$ and $y$ is an element of the unipotent family $C_{u_{0}^{\vee}}$ parametrized by $u_{0}^{\vee}$. The element $y$ will not be important for us. Let $\widetilde{C}_{u_{0}^{\vee}}=C_{\widetilde{u}_{0}^{\vee}}$ be the family obtained by applying Alvis-Curtis duality to $C_{u_{0}^{\vee}}$. Let $d_{S}^{\vee}$ be the duality map in this finite field setting, this is defined exactly as above, see [Lu1, §13.3].
Theorem 3.2 ([Lu2, Theorem 11.2]).
@@
${}^{K}\!\underline{\mathsf{WF}}(\pi(x,\sigma))=(d_{S}^{\vee}(\tau u_{0}^{\vee}),x).$ (3.1)
Since we are in a depth-zero case, the Langlands parameter $\varphi$ is trivial on the wild inertia group, and the image $\varphi(I_{\mathsf{k}})$ is a cyclic group generated by an element $\lambda\in{}^{L}\!G$ of finite order prime to $q$. Hence ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\lambda}$ is a pseudo-Levi subalgebra of $\mathfrak{g}^{\vee}$. Conjecture 1.6 is equivalent to
@@ 3.1. Regular supercuspidal representations
For all regular depth-zero supercuspidal representations [KV, DBR], the conjecture is trivially true since in that case ${\mathfrak{h}}^{\vee}$ is just a Cartan subalgebra of $\mathfrak{g}^{\vee}$, so the only possible Langlands nilpotent element is 00. Hence the conjecture just says that ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq{{\mathcal{O}}}_{p}$, where ${{\mathcal{O}}}_{p}$ is the principal ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit, which is true, but trivial. On the other hand, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={{\mathcal{O}}}_{p}$ if and only if $\pi$ is a generic representation, and this is not always the case. In general, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ can be equal to the ${\mathbf{G}}(\bar{\mathsf{k}})$-saturation of the principal nilpotent orbit in a maximal pseudo-Levi, as we explain next.
@@
By abuse of notation, denote $x$ also the preimage in $X$ of an element in $X/(1-w\theta)X$. Let $\pi(\lambda,w,0,x)$ denote the supercuspidal representation attached to $(\lambda,w,x)$ in loc. cit. The 00 here stands for the zero nilpotent orbit, i.e., the image of the $SL(2)$ under the Langlands parameter must be trivial (since ${\mathfrak{g}^{\vee}}^{\lambda(I_{\mathsf{k}})}=\mathfrak{t}^{\vee}$).
@@
The claim follows at once, after we recall the construction of $\pi(\lambda,w,0,x)$. The element $x$ defines a cocycle $u_{x}\in H^{1}({\mathsf{k}},{\mathbf{G}})$ which gives the inner twist $\omega_{x}$ of ${\mathbf{G}}$ and $F_{x}$ a twisted Frobenius map and a maximal parahoric subgroup ${\mathbf{G}}_{x}$. Then ${\mathbf{G}}_{x}^{F_{x}}$ is a maximal parahoric subgroup of ${\mathbf{G}}^{\omega_{x}}({\mathsf{k}})$. The pair $(w,x)$ defines an $F_{x}$-minisotropic torus $T_{x}$ and $\lambda$ defines a depth-zero character $\chi_{x}$ of $T_{x}$. Then Deligne-Lusztig induction defines an irreducible cuspidal representation of the reductive quotient $\overline{{\mathbf{G}}_{x}^{F_{x}}}$
@@
4. Depth-zero representations of classical groups
In [LuSt], Lust and Stevens constructed L-packets of depth-zero supercuspidal representations of classical groups. We will verify Conjecture 1.6 for these $L$-packets. Let ${\mathcal{G}}$ be a connected reductive group of classical type with connected center over ${\mathbb{F}}_{q}$, $q$ odd. Let $F$ be the Frobenius map. Let ${\mathcal{G}}^{*}$ be the dual group over ${\mathbb{F}}_{q}$ with dual Frobenius also denoted $F$. Let $\mathsf{Irr}{\mathcal{G}}^{F}$ be the set of equivalence classes of irreducible complex ${\mathcal{G}}^{F}$-representations. According to Lusztig’s classification, we can partition